Suppose many capacitors, each with 𝐢 =90.0 μ𝐹, are connected in parallel across a battery

with a potential difference of βˆ†π‘‰ =160.0 𝑉. How many capacitors are needed to store 95.6 J

of energy?


1
Expert's answer
2021-11-08T17:12:26-0500

"C=90\mu F\\∆V=160volt\\U=95.6J"

"U=\frac{1}{2}C'∆V^2"

Assume n capacitor connected in parallel combination

C'=nC

"U=\frac{1}{2}nC∆V^2"


"n=\frac{2U}{C∆V^2}=\frac{2\times95.6\times}{90\times10^{-6}\times160\times160}=82"

n=82